# Exact staggered dimer ground state and its stability in a two-dimensional magnet

### v2

@@ I Introduction
The importance of magnetism in matter got renewed attention after the breakthrough and remarkable discovery of high-$T_{c}$ superconductors. The first reported high-$T_{c}$ superconductor was synthesized by optimal doping of Ba in the material La2CuO4 [1], wherein the Cu2+ ions form magnetically weakly coupled layers with an underlying square lattice. The undoped lanthanum copper oxide is an antiferromagnetic insulator below the Néel temperature ($T_{N}\sim 325$ K). It goes into a superconducting state (up to the critical temperature $T_{c}\sim 35.1$ K) when sufficient holes are generated through the replacement of La3+ ions by the Ba2+ ions. The physical properties of these materials are significantly different from the zero-temperature superconductors, so the Bardeen–Cooper–Schrieffer theory [2] does not work in these cases. Anderson attempted this intriguing problem, and to understand the high-$T_{c}$ superconductivity, he proposed the resonating valence bond (RVB) theory [3]. This theory uses the highly degenerate valence bond states [4, 5]. Here, the spin-singlet pairs of the insulating phase become charged superconducting pairs upon adequate doping. This theory has had substantial success, but the search for a more complete theory is still ongoing. Nevertheless, most researchers believe that magnetism plays an indispensable role in the high-$T_{c}$ superconductors [6, 7].
Interacting spin systems with the valence-bond ordered ground states have been in focus for many decades due to their theoretical and technological importance. A straightforward prototype spin system with competing interactions is the Majumdar-Ghosh (MG) model [8, *Majumdar1969b]. It is a spin-$1/2$ Heisenberg model on a one-dimensional chain lattice with nearest- and next-nearest-neighbor antiferromagnetic exchange interactions $J_{1}$ and $J_{2}$, respectively. Its ground state energy at $J_{2}/J_{1}=1/2$ (MG point) is doubly degenerate, and the two nearly orthogonal dimer states span the ground state manifold. As one increases the interaction ratio $J_{2}/J_{1}$, this model undergoes a quantum phase transition from a gapless quasi-long-range ordered state to a gapped dimer state at a quantum critical point, $J_{2}/J_{1}\sim 0.24$ [10, 11]. Many quasi-one-dimensional materials show the signature of having the characteristic features of the MG model. The multiferroic compound CuCrO4 finds its place close to the MG point [12], and the zigzag antiferromagnet (N2H5)CuCl3 seems to be at the critical point [13].
@@
There is another class of Hamiltonians in which frustration (as in MG and SS models) and spatial anisotropy play essential roles. One such example is the Nersesyan-Tsvelik (NT) model [30], which transforms to an isotropic $J_{1}$-$J_{2}$ Heisenberg antiferromagnet [31] in the absence of anisotropy. The NT model is a frustrated and spatially anisotropic spin-$1/2$ Heisenberg antiferromagnet on a square lattice. Here, the spatial anisotropy introduces, in general, unequal exchange interactions $J_{1}^{\parallel}$ and $J_{1}^{\perp}$ on the horizontal and vertical nearest-neighbor bonds, respectively. The competing next-neighbor exchange interaction $J_{2}$ induces frustration in the system. In this model, an RVB phase with deconfined spinons emerges along the line $J_{1}^{\perp}=2J_{2}$ for $J_{1}^{\perp},J_{2}\ll J_{1}^{\parallel}$  [30]. As one goes beyond the limit $J_{1}^{\perp}/J_{1}^{\parallel}\ll 1$, a columnar dimer phase appears around the line $J_{1}^{\perp}=2J_{2}$ with increasing domain in the phase diagram  [32, 33]. The compound $(\text{NO})[\text{Cu}{({\text{NO}}_{3})}_{3}]$ appears a potential candidate for an experimental realization of the NT model [34].
@@ II Model
where the summation runs over all possible triangles with vertices (or sites) $i,j$, and $k$ such that each triangle contains one dimer, one non-dimer nearest-neighbor bond, and one next-nearest-neighbor bond. If there are $N$ total dimers in the lattice with periodic boundary conditions, the total triangles will be $4N$. Equivalently, it means that each dimer contributes four triangles (see Fig. 1). The spin operators $\mathbf{S}_{i},\mathbf{S}_{j}$, and $\mathbf{S}_{k}$ present at the vertices of a triangle define the projector operator $\mathcal{P}_{3/2}(i,j,k)$ as follows:
@@
Figure 5: The averaged spin-spin correlation on the three types of bonds for the $4\times 7$ cluster is shown here. The error bar is negligible at the exact point but significantly increases on dimers as one goes farther away from $J_{2}/J_{1}=0.5$.
@@
It would be interesting to compare the model (1) with the MG and SS models. All these models involve only nearest-neighbor and next-nearest-neighbor exchange interactions and have dimer ground states at specific points in the coupling space. However, the SS ground state in the SS model exists in a broad range. Except for MG, the ground states are unique. Our and SS models have a spatial anisotropy in the nearest-neighbor and next-nearest-neighbor interactions, respectively. However, the MG model is isotropic in both interactions. Structurally, the SS and our models correspond to the bond-depleted $J_{1}$-$J_{2}$ Heisenberg models on a square lattice. One wipes out $3/4$th next-nearest-neighbor bonds in the SS case, while we eradicate $1/4$th nearest-neighbor bonds in our spatially anisotropic (or dimerized) model.
@@ V Conclusions
The above model can also be extended in three dimensions but with a different solvable point for an exact staggered dimer ground state. We show such a possibility on the left in Fig. 12. Here, the two-dimensional staggered dimer lattices of the type shown in Fig. 1 are stacked directly above one another along the $z$-axis such that all even layers (imagine infinite in extent) are translated by the same $a\hat{y}$ amount. The lattice planes normal to the $y$- and $z$-axes have dimers. However, the lattice planes perpendicular to the $x$-axis are without dimers and form an isotropic $J_{1}$-$J_{2}$ Heisenberg model on square lattices. We expect that this proposed model on the three-dimensional lattice with exchange interactions $J_{D}$, $J_{1}$, $J_{2}$ will have a staggered dimer ground state when one sets $J_{D}=4J_{1}$ and $J_{2}/J_{1}=1/2$. It is because the number of triangles that share the same dimer becomes double the corresponding value in the two-dimensional model 1, while nothing changes for non-dimer bonds.

### v3

@@ I Introduction
Frustrated magnetism is a fascinating and vigorous field of research that attracts theoreticians, experimentalists, and material scientists equally due to emerging new understandings and challenges in matters with competing interactions [1, 2, 3, 4, 5, 6, 7, 8, 9]. Interacting spin systems with the valence-bond ordered ground states is one subclass of this field, which has been in focus for many years due to their novel physics about phases and their transitions. A straightforward prototype spin system with competing interactions is the Majumdar-Ghosh (MG) model [10, *Majumdar1969b]. It is a spin-$1/2$ Heisenberg model on a one-dimensional chain lattice with nearest- and next-nearest-neighbor antiferromagnetic exchange interactions $J_{1}$ and $J_{2}$, respectively. Its ground state energy at $J_{2}/J_{1}=1/2$ (MG point) is doubly degenerate, and the two nearly orthogonal dimer states span the ground state manifold. As one increases the interaction ratio $J_{2}/J_{1}$, this model undergoes a quantum phase transition from a gapless quasi-long-range ordered state to a gapped dimer state at a quantum critical point, $J_{2}/J_{1}\sim 0.24$ [12, 13]. Many quasi-one-dimensional materials show the signature of having the characteristic features of the MG model. The multiferroic compound CuCrO4 finds its place close to the MG point [14], and the zigzag antiferromagnet (N2H5)CuCl3 seems to be at the critical point [15].
@@
There is another class of Hamiltonians in which frustration (as in MG and SS models) and spatial anisotropy play essential roles. One such example is the Nersesyan-Tsvelik (NT) model [32], which transforms to an isotropic $J_{1}$-$J_{2}$ Heisenberg antiferromagnet [33] in the absence of anisotropy. The NT model consists of a set of identical spin chains that are arranged horizontally one after the other in two-dimensional space. Each spin chain carries exchange coupling $J_{1}^{\parallel}$ for any pair of neighboring spins of the chain and interacts with its nearest-neighboring spin chains through transverse and diagonal exchange couplings, denoted by $J_{1}^{\perp}$ and $J_{2}$, respectively. In this model, all couplings are antiferromagnetic and satisfy the condition $J_{1}^{\perp},J_{2}\ll J_{1}^{\parallel}$. The line $J_{1}^{\perp}=2J_{2}$ classically represents a first-order phase transition between the Néel and Stripe magnetic orderings. However, the analytical and numerical studies encapsulating the quantum nature of spins reveal a dimer phase between the Néel and Stripe orderings, and the dimer phase undergoes first-order transitions with both magnetically ordered phases [34, 35]. The compound $(\text{NO})[\text{Cu}{({\text{NO}}_{3})}_{3}]$ appears a potential candidate for an experimental realization of the NT model [36].
@@ II Model
where the summation runs over all possible triangles with vertices (or sites) $i,j$, and $k$ such that each triangle contains one dimer, one non-dimer nearest-neighbor bond, and one next-nearest-neighbor bond. If there are $N$ total dimers in the lattice with periodic boundary conditions, the total triangles will be $4N$. Equivalently, it means that each dimer contributes four triangles (see Fig. 1). The spin operators $\mathbf{S}_{i},\mathbf{S}_{j}$, and $\mathbf{S}_{k}$ present at the vertices of a triangle define the projection operator $\mathcal{P}_{3/2}(i,j,k)$ as follows:
@@
Figure 5: The averaged spin-spin correlation on the three types of bonds for the $4\times 7$ cluster with open boundaries along the vertical direction is shown here. The error bar is negligible at the exact point but significantly increases on dimers as one goes farther away from $J_{2}/J_{1}=0.5$.
@@
It would be interesting to compare the model (1) with the MG and SS models. All these models involve only nearest-neighbor and next-nearest-neighbor exchange interactions and have dimer ground states at specific points or in a range in the coupling space. Except for the MG model, the dimer ground states are unique. Our and SS models have a spatial anisotropy in the nearest-neighbor and next-nearest-neighbor interactions, respectively. However, the MG model is isotropic in both interactions. Structurally, the SS and our models correspond to the bond-depleted $J_{1}$-$J_{2}$ Heisenberg models on a square lattice. One wipes out $3/4$th next-nearest-neighbor bonds in the SS case, while we eradicate $1/4$th nearest-neighbor bonds in our spatially anisotropic (or dimerized) model. Moreover, our model also shows the staggered dimer ground state for any $J_{D}\geq 2J_{1}$ if one sets the ratio $J_{2}/J_{1}$ equal to $1/2$, as the SS ground state in the SS model exists in a broad range. This can be understood as follows. At $J_{2}/J_{1}=1/2$, using Eq. (3), the Hamiltonian (1) can be written as
$H=E_{c}+\frac{J_{1}}{2}\sum_{\{(i,j,k)\}}\left[\kappa\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}+\frac{3}{2}\,\mathcal{P}_{3/2}(i,j,k)\right],$ (5)
where $E_{c}=-(3/2)J_{1}N$, $\kappa=\left(J_{D}/2J_{1}-1\right)$, and the indices $i,j$ are endpoints of a dimer. The product state (4) is an eigenstate of the Hamiltonian (5) with the eigenenergy $-(3/4)J_{D}N$, independent of the coupling $J_{1}$. As we know the slack inequalities $-3/4\leq\expectationvalue{\mathbf{S}_{i}\cdot\mathbf{S}_{j}}\leq 1/4$ and $0\leq\expectationvalue{\mathcal{P}_{3/2}(i,j,k)}\leq 1$ of expectation values hold for any arbitrary state that is a linear combination of the energy eigenstates, the staggered dimer state will be the ground state for any $\kappa\geq 0$ (see Fig. 6).
Figure 6: This figure displays the ground state energy per dimer (in units of $J_{D}$) calculated from the exact diagonalization and analytical calculations; here, we keep the exchange constant $J_{1}$ to unity. The diagonalization data is produced for the $6\times 4$ lattice with fully- and semi-periodic boundary conditions; here, semi-periodic means boundaries along the vertical direction are open.
It turns out that the dimer ground state also persists for a range with $\kappa<0$, similar to the SS model [17]. We arrive at the same conclusions of the Hamiltonian with fully-periodic boundary conditions even if the boundaries along the vertical directions are kept open, except that the range of $\kappa$ for the dimer state being a ground state is slightly wider in the semi-periodic case.
@@ V Conclusions
The above model can also be extended in three dimensions but with a different solvable point for an exact staggered dimer ground state. We show such a possibility on the left in Fig. 13. Here, the two-dimensional staggered dimer lattices of the type shown in Fig. 1 are stacked directly above one another along the $z$-axis such that all even layers (imagine infinite in extent) are translated by the same $a\hat{y}$ amount. The lattice planes normal to the $y$- and $z$-axes have dimers. However, the lattice planes perpendicular to the $x$-axis are without dimers and form an isotropic Heisenberg model with only nearest-neighbor interactions on square lattices. We expect that this proposed model on the three-dimensional lattice with exchange interactions $J_{D}$, $J_{1}$, $J_{2}$ will have a staggered dimer ground state when one sets $J_{D}=4J_{1}$ and $J_{2}/J_{1}=1/2$. It is because the number of triangles that share the same dimer becomes double the corresponding value in the two-dimensional model 1, while nothing changes for non-dimer bonds.